OpenStudy (khushi):

F (x) = {a+b, a+bx4 5 When x=2 x= <2, >2, =2

OpenStudy (khushi):

Please explain it

OpenStudy (khushi):

a+bx+ 4

OpenStudy (will.h):

Can you try to post the question in a better format can't understand anything out of this

OpenStudy (khushi):

F(x)= {a+b, a+bx+4 5 x<2 x=2 x>2 Find value of a and b if the function f given is x=2

OpenStudy (phi):

It looks like you have a "piecewise" function definition see http://www.mathsisfun.com/sets/functions-piecewise.html for examples.

OpenStudy (phi):

to find a and b we need more info. For example, do they say the function is "continuous" ? do they give values for f(0) or f(2) (for example)

OpenStudy (khushi):

Yes function is continuous at x=2

OpenStudy (khushi):

No value is given it starts with f (x)= { like this...

OpenStudy (phi):

if the function is continuous, then as x->2 the top line a+b must get close to f(2) the middle line a+bx+4 is f(2) (or because x=2) a+2b+4 = f(2) and the last line (for x>2) should also "approach" f(2) as x gets close to 2 (in other words, "near" 2, all three definitions should be the same number)

OpenStudy (phi):

because the last line says F( 2+(tiny number) )= 5 we expect the middle line a+2b+4 = 5 and a+b = 5 that is enough info to solve for a and b can you do that?

OpenStudy (khushi):

Can u please help? Its an assignment given to be completed in very short time and m not from math field so getting problem.. Please help if u can